Show that the sample mean x, is sufficient for estimating the parameter λof the Poisson distribution.
1
Expert's answer
2020-03-09T14:09:44-0400

We will be using Neymann factorization Theorem for proving this. Let us first write down the joint density of independent Poisson random variables with the same parameter.


"f_{\lambda}(X_1,\ \ldots, X_n) = e^{-n\lambda}\times\frac{\lambda^{\sum X_i}}{\prod_i X_i!}"

This can be written as "f_{\lambda}(X_1,\ \ldots, X_n) = h(X_1,\ \ldots, X_n) \cdot g(\lambda, T(X))"


Here, T(X) = "\overline{X}" by choosing "h(X_1,\ \ldots, X_n) = \frac{1}{\prod_i X_i!}" and "g(\lambda, T(X)) = e^{-n\lambda}\times\lambda^{n\overline{X}}"


Now, Neymann factorization states that if we can decompose the density in the above way then T(X) which is the sample mean is sufficient for estimating the parameter "\lambda" . Hence proved.